34.A rectangle has a perimeter of 24 units if the width is 4 units, which is the length?
Answer by checkley79(3341)

P=2L+2W
24=2L+2*4
24=2L+8
2L=24-8
2L=16
L=16/2
L=8 UNITS (ANSWER).
PROOF:
24=2*8+2*4
24=16+8
24=24